OpenStudy (anonymous):

I think its B, I will give a medal if your right? Solve the equation: 4x2 = 36 {6, -6} {3, -3} {2/3, -2/3} {3/2, -3/2}

Directrix (directrix):

I think the answer may be "B." Let's see.

Directrix (directrix):

4x2 = 36 (1/4) (4x^2) = (1/4)(36) x^2 = 9 x = 3 or -3 I agree. @GreenBeanDemphsey

OpenStudy (anonymous):

thank you!

Directrix (directrix):

Happy to help.